My Algebra final is today, I need to complete 2 word problems, 1) A credit union loaned out $50000, part at an annual rate of 5% ad the rest at an annual rate of 10%. The collected combined simple interest was $3000 that year. How much was loaned out at 5%? I have attempted to us formula of I=Prt, but I am unable to figure out were to start, How much of the year is part, Please help med set it up and completely solve this problem. Second problems is: A New elementary algebra textbook sells for $80 and a used elementary algebra textbook sells for $60. How many new textbooks are sold if the receipts for 125 new and used textbooks is $9060? I really do need help. I been working on this for 1 whole day. Thank You Kathy

Answer by KMST(5328)

1) There were 2 loans accruing interest for the whole year. The amount loaned at at 5%, in $, was x.
The amount of the other loan, in $, was 50000 - x, and that was loaned at 10%.
The interest on the first loan was 0.05x,
because the rate was r = 5% = 5/100 = 0.05,
The principal, P, was x, and the time, t, was 1 year.
You can calculate the interest of the other loan the same way,
taking into account that for that loan
P= 50000 - x , and r = 10% = 10/100 = 0.1 .
The interest from both loans, in $, adds up to 3000.
From that, you make your equation.
2) x = number of new textbooks sold,
So the total paid in receipts for those new textbooks, in $, was 80x, and
125 - x = number of used textbooks sold.
You can figure out an expression for how much was paid for the used textbooks,
and knowing that the total amounts (in$) paid for new and used textbooks add up to
9060, you can set up your equation to find x.

If that is not enough help, you need an algebra course that really helps you understand what algebra is and how it relates to word problems simulating real life. It would not really help if I solved the problems for you, and you passed the test without really understanding how those problemas should be solved.